The femoral artery is the large artery that carries blood to the leg. A person's femoral artery has an inner diameter of 1.3 cm. What is the resistance of a 19-cm-long column of blood in this artery?

     ______________?
Answer #1

resistance = resistivity * length /area = resistivity * 19/(pi*1.3*1.3/4) = resistivity * 14.31
Answer #2

Rosrstivi = Td2-( nxCooer= o.oto12 m2 し,う 4 O 0D013